OpenStudy (anonymous):

(___)^2=(secx-1)(secx+1) fill in the blank

OpenStudy (raden):

(sec(x)+1)(sec(x)−1)=sec^2 (x)−1 = tan^2 (x)

OpenStudy (raden):

so, tan(x) is the answer

OpenStudy (anonymous):

awesome thank you

OpenStudy (raden):

welcome